using the following table, find f(x) as a polynomial in x:

x: – 1 0 3 6 7

f(x): 3 – 6 39 822 1611.


1
Expert's answer
2021-01-05T14:40:18-0500

Let "f(x)=ax^4+bx^3+cx^2+dx+e", then we have a system of equations

"\begin{cases}
f(-1)=a-b+c-d+e=3\\
f(3)=81a+27b+9c+3d+e=39\\
f(6)=1296a+216b+36c+6d+e=822\\
f(7)=2401a+343b+49c+7d+e=1611\\
f(0)=e=-6
\end{cases}"

Extended matrix of the system is

"\begin{pmatrix}
1&-1&1&-1&1&3\\
81&27&9&3&1&39\\
1296&216&36&6&1&822\\
2401&343&49&7&1&1611\\
0&0&0&0&1&-6
\end{pmatrix}"

Solve the system by Jordan method

"\begin{pmatrix}
1&-1&1&-1&1&3\\
81&27&9&3&1&39\\
1296&216&36&6&1&822\\
2401&343&49&7&1&1611\\
0&0&0&0&1&-6
\end{pmatrix}\rightarrow"

"\rightarrow\begin{pmatrix}
1&-1&1&-1&0&9\\
81&27&9&3&0&45\\
1296&216&36&6&0&828\\
2401&343&49&7&0&1617\\
0&0&0&0&1&-6
\end{pmatrix}\rightarrow"

"\rightarrow\begin{pmatrix}
1&-1&1&-1&0&9\\
0&108&-72&84&0&-684\\
0&1512&-1260&1302&0&-10836\\
0&2744&-2352&2408&0&-19992\\
0&0&0&0&1&-6
\end{pmatrix}\rightarrow"

"\rightarrow\begin{pmatrix}
1&-1&1&-1&0&9\\
0&9&-6&7&0&-57\\
0&36&-30&31&0&-258\\
0&49&-42&43&0&-357\\
0&0&0&0&1&-6
\end{pmatrix}\rightarrow"

"\rightarrow\begin{pmatrix}
1&0&\frac{1}{3}&-\frac{2}{9}&0&\frac{8}{3}\\
0&9&-6&7&0&-57\\
0&0&-6&3&0&-30\\
0&0&-\frac{28}{3}&\frac{44}{9}&0&-\frac{140}{3}\\
0&0&0&0&1&-6
\end{pmatrix}\rightarrow"

"\rightarrow\begin{pmatrix}
1&0&\frac{1}{3}&-\frac{2}{9}&0&\frac{8}{3}\\
0&9&-6&7&0&-57\\
0&0&-2&1&0&-10\\
0&0&-84&44&0&-420\\
0&0&0&0&1&-6
\end{pmatrix}\rightarrow"

"\rightarrow\begin{pmatrix}
1&0&0&-\frac{1}{18}&0&1\\
0&9&0&4&0&-27\\
0&0&-2&1&0&-10\\
0&0&0&2&0&0\\
0&0&0&0&1&-6
\end{pmatrix}\rightarrow"

"\rightarrow\begin{pmatrix}
1&0&0&0&0&1\\
0&9&0&0&0&-27\\
0&0&-2&0&0&-10\\
0&0&0&1&0&0\\
0&0&0&0&1&-6
\end{pmatrix}\rightarrow"

"\rightarrow\begin{pmatrix}
1&0&0&0&0&1\\
0&1&0&0&0&-3\\
0&0&1&0&0&5\\
0&0&0&1&0&0\\
0&0&0&0&1&-6
\end{pmatrix}"

We obtain "a=1, b=-3, c=5, d=0, e=-6", that is "f(x)=x^4-3x^3+5x^2-6"

Answer: "f(x)=x^4-3x^3+5x^2-6"